OpenStudy (anonymous):

Simplify the expression 4 over quantity of 3 x to the negative second power.

OpenStudy (mathstudent55):

Is this the porblem? \( \dfrac{4}{(3x)^{-2} } \)

OpenStudy (anonymous):

yes!

OpenStudy (mathstudent55):

Look at the following rule of exponents: \(a ^{-n} = \dfrac{1}{a^n} \)

OpenStudy (mathstudent55):

What the rule tells you is that if you have a negative exponent, it's the same as having a positive exponent in the denominator. Now let's use that rule of exponents for a negative exponent in the denominator to see what happens to the negative exponent.

OpenStudy (mathstudent55):

\( \dfrac{1}{a^{-n}} \) \(= \dfrac{~~~1~~~}{\dfrac{1}{a^n}} \) \(= 1 \div \dfrac{1}{a^n} \) \(= 1 \times a^n \) \( = a^n\)

OpenStudy (mathstudent55):

From here you see that a negative exponent in the denominator is a positive exponent in the numerator.

OpenStudy (mathstudent55):

Now let's look at your problem. \(\dfrac{4}{(3x)^{-2} }\) \(= 4 \times (3x)^{2} \) Now can you finish it?

OpenStudy (anonymous):

\[36x^{2}?\]

OpenStudy (mathstudent55):

\(= 4 \times (3x)^{2}\) \(= 4 \times 9x^2\) \( = 36x^2\) You are correct. Good job!

OpenStudy (anonymous):

thank you!!